OpenStudy (anonymous):

erica bought 3 1/2 yards of fabric. if she uses 2/3 of the fabric to make a curtain, how much will she have left?

OpenStudy (anonymous):

\[\huge 3\frac 12 - \frac 23\] solve that and you have your answer

OpenStudy (anonymous):

i got 2 5/6 but on my prep sheet it says the awnser is 1 1/6

OpenStudy (anonymous):

it isn't 2/3 but 2/3 OF the 3.5

OpenStudy (anonymous):

first make 3 1/2 into improper fraction \[\huge 3 \frac 12 \implies \frac 72\]

OpenStudy (anonymous):

ye i know tht part

OpenStudy (anonymous):

oh yes.

OpenStudy (anonymous):

didn't see that

OpenStudy (anonymous):

\[\huge 3 \frac 12 - \frac 23 (3\frac 12)\] that would be the right formula

OpenStudy (anonymous):

3 1/2-14/5?

OpenStudy (anonymous):

so you should have \[\huge \frac 72 - \frac 23 (\frac 72)\] so that's \[\huge \frac 72 - \frac {14}6\] then \[\huge \frac{21}6 - \frac{14}6\]

OpenStudy (anonymous):

lol whoops

OpenStudy (anonymous):

that should result to 1 1/6

OpenStudy (anonymous):

but the thing is wut made is to different from being a regualr subtraction problem

OpenStudy (anonymous):

that would be the one that confused me too..the line "he uses 2/3 of the fabric" so you have to get 2/3 of the fabric first before subtracting

OpenStudy (anonymous):

OOOOOOOOOOOOOOOOOOOOOOOOOO

OpenStudy (anonymous):

of the total